# Efficient Iterative Proximal Variational Inference Motion Planning

### v3

@@ 4.1 Sparse Motion Planning Factor Graph
which can be broken apart into the calculation of the expectations of $\psi$
@@
The motion planning factor graph is shown in the following Fig. 1. It contains two types of factors, the $f_{i,i+1}(x_{i},x_{i+1})$ represents the prior factor enforcing the kinodynamics constraints, and the $f_{i}(x_{i})$ are the collision factors that encourage the trajectory remain within obstacle-free regions.
@@
This factorized structure can be leveraged to distribute the computation of $\psi_{d}$ on a GPU. Given the factor graph structure, the collision costs $\psi_{d}=\lVert h\left(x\right)\rVert^{2}_{\Sigma_{\rm obs}}$ can be factorized into
@@
where the factorized potentials are defined as the collision cost on each individual state on the trajectory
$\psi^{d}_{i}$ $\triangleq\lVert h_{i}(q_{i})\rVert_{\sigma_{\rm obs}}^{i}.$
The factor-level marginal Gaussian variables are linearly mapped from the joint Gaussian
@@
Nest, we explore these sparsity patterns in the proposed Gaussian distributions.
@@ 4.2 Marginal updates
By Stein’s Lemma, we write (20a) and (20b) as
@@
Since we can decompose the $\psi(\mathbf{x})$ to marginal level
@@
where $\mathbf{x}_{i}=M_{i}\mathbf{x}$ is a subset of the variables relevant to the $i$th factor as defined in (23). Applying the factorization (25) to (24a), we obtain
@@
For the $i$-th factor, we can simplify the gradient and expectation to being over just $\mathbf{x}_{i}$, since $\psi_{i}$ only depends on $\mathbf{x}_{i}$. We have a similar result for (24b)
@@
Apply Stein’s Lemma again to (26) and (27), we have
@@ 4.3 Sparse Quadratures for Expectation Estimations
Our primary goal is to estimate the expectation (28) on the marginal levels. Gauss-Hermite quadrature approximations are widely used in filtering literatures [50, 32], as well as motion planning [14, 15]. Cubature approximations are accurate and efficient for lower dimensional problems. However, for higher dimensional problems, a straight-forward tensor product of the full-grid sigma points scale exponentially with the dimension, making it infeasible for high-dimensional problems [51].
Sparse-grid quadrature rules (Smoyak’s rules) [52, 53] ignore the cross terms among different dimensions to mitigate the curse of dimensionality, resulting in a quadrature rule with polynomial dimensional dependence. In this work, we adopt the sparse-grid Gauss-Hermite quadrature rules to estimate (28). To integrate a scalar function $\varphi(\mathbf{x})$
[…]
@@
    Compute the $p$ roots of a $p\;th$ order Hermite polynomial, also denoted as sigma points: $\xi=[\xi_{1},\dots,\xi_{p}].$
@@
[…]
Lemma 1 (Smoyak sparse quadrature complexity).
For a $n-$variate function, the number of computations needed by a sparse grid quadrature that is exact for $p=2k_{q}-1$ total polynomial order is bounded by [53]
@@
Algorithm 1 PGVIMP (LTV system).
@@ 5      Perform P-GVIMP (Algorithm 1) on the linearized system
6      $\left\{\left(\bar{z}^{i+1}_{t},\bar{\Sigma}^{i+1}_{t}\right)\right\}\leftarrow$ P-GVIMP$\left(A^{i}_{t},a^{i}_{t},B^{i}_{t}\right)$
@@
Algorithm 2 i-PGVIMP (Nonlinear system).
5 Parallel GVIMP For Linear Time-Varying Systems
We first introduce a Parallel GVIMP algorithm on GPU for a linear time-varying (LTV) system.
@@ 5.1 Belief propagation for marginal covariances update
Computing the marginal updates (26) and (27) requires computing the marginal mean and covariance of the Gaussian distributions $q_{i}$. A major bottleneck is computing the marginal covariances. A brute force inverse computation of the joint covariance from $\Sigma^{-1}_{\theta}$ has the complexity $O((N\times n)^{3})$, which is a cubic function in $N\gg n$ in motion planning.
Leveraging the sparse factor graph, this section introduces the Gaussian Belief Propagation (GBP) [27, 28] to compute the marginals over the factor graph. Gaussian distribution $q\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})$ can be written in the canonical form
$q(x)\propto\exp(-\frac{1}{2}x^{\top}\Lambda_{\theta}x+\eta_{\theta}^{\top}x)$ (30)
where $\Lambda_{\theta}=\Sigma_{\theta}^{-1}$ denotes the precision matrix, and $\eta_{\theta}=\Sigma_{\theta}^{-1}\mu_{\theta}$ denotes the information vector
To simplify the computation, we can assume a shifted Gaussian distribution $r\sim\mathcal{N}(0,\Sigma_{\theta})$, since we only want to obtain the marginal covariance. Then $r(x)$ can be written as
@@
The precision matrix $\Lambda_{\theta}$ has such sparsity pattern
$\Lambda_{\theta}\!=\!\begin{bmatrix}\Lambda_{00}\!\!&\Lambda_{01}\!\!&&&&\\ \Lambda_{10}\!\!&\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &&&\dots&&\\ &&&&\!\!\Lambda_{(N-1)(N-1)}\!\!&\Lambda_{(N-1)N}\\ &&&&\!\!\Lambda_{N(N-1)}\!\!&\Lambda_{NN}\end{bmatrix},$ (32)
and the Gaussian distribution (31) can be factorized into
$r(x)\propto\prod_{i=0}^{N}f_{i}(x_{i})\ \prod_{i=0}^{N-1}f_{i}(x_{i},x_{i+1})$ (33)
with
@@
Gaussian Belief Propagation is an algorithm for calculating the marginals of a joint distribution via local message passing between nodes in a factor graph [28]. Since all of these factors are Gaussian, we can use this method directly.
Message passing on the factor graph can be divided into two types. The message passed from variables to factors, denoted as $m_{x_{i}\rightarrow f_{j}}$, and the message passed from factors to variables, denoted as $m_{f_{j}\rightarrow x_{i}}$. Messages can be obtained by
@@
After computing all the messages in the factor graph, the beliefs of variables can be obtained by taking the product of incoming messages:
$b_{i}\left(x_{i}\right)=\prod_{s\in N(i)}m_{f_{s}\rightarrow x_{i}}$ (35)
since the messages here are also Gaussian distribution, we can obtain the belief parameters $\Lambda_{bi}$ by
@@
The time complexity of inverting a $n\times n$ matrix is $O(n^{3})$, leading to a total complexity $O(N\times n^{3})$ for GBP.
@@ 5.2 Parallel Collision Checking on Dense Trajectory
One of the most computationally heavy modules for motion planning algorithms is the collision-checking module on a dense trajectory [4, 11]. This is also the case for variational motion planning algorithms [14, 15], where the collision-checking is required in a probabilistic sense, i.e., the goal is to compute the expected collision costs with respect to the proposal distribution, which is more challenging than computing a deterministic collision cost.
In our proposed method, we compute the expected collision costs by distributing the computation to each factor in the factor graph, as introduced in Section 4.2. Then, at marginal levels, we deploy the sparse GH quadratures to compute the marginal collision costs, as introduced in Section 4.3. After specifying the desired precision of this approximation, the quadrature method computes the sigma points and the weights, which are fixed throughout the experiment, and the expectation approximation is done by computing the matrix operations (29). Thus, one straightforward way of distributing the expectation approximation is by distributing the matrix multiplication (29), which is implemented in our proposed framework. The graphical demonstration of the parallel collision checking is shown in Fig. 2.
@@
5.3 PGVIMP Algorithm and Complexity Analysis
The PGVIMP algorithm is summarized in the following Algorithm 1. This algorithm serves as a sub-rouitne in motion planning for nonlinear systems by leveraging iterative linearization.
Complexity Analysis. We now analyze the time complexity of the proposed parallel GVIMP algorithm. The state dimension is $n$, and time discretization is $N+1$. The factor graph has $L=2N+3$ factors, the maximum dimension being $d=2n$. A serial computation of the nonlinear factors leads to a linear dependence on $N$, and the total complexity is $O(N\times n^{k_{q}})$, where $k_{q}$ is the desired precision in the GH-quadrature [15]. Experimentally, to obtain a precise estimate of the collision costs, $k_{q}\geq 6$ is required. The parallel nonlinear factor computation is of complexity $O(n)$, and the complexity of computing the marginal covariances for the tree-structured factor graph using GBP is of $O(N\times n^{3})$. The overall complexity of the algorithm is therefore reduced from $O(N\times n^{k_{q}})$ to $O(N\times n^{3})$, which is not dependent on the quadrature precision $k_{q}$.
@@ 6 Iterative P-GVIMP for Nonlinear Systems
Many robot dynamical systems are nonlinear. The P-GVIMP algorithm described in the previous sections constrained the system to be LTV. To handle the nonlinearity, this section introduces our proposed Iterative Parallel GVIMP Algorithm (i-PGVIMP).
@@ 6.1 Gaussian Assumptions for Nonlinear Transition
It is in general difficult to solve the FPK for arbitary nonlinear stochastic system (1). Gaussian assumption is popular in the robotics applications, especially in the filtering and estimation [32, 31]. It assumes the state distribution follows
@@
where $\mathbb{E}_{\mathcal{N}}\left[\psi(\cdot)\right]\triangleq\int\psi(x)\mathcal{N}(x)$ denotes the expected values under Gaussian distributoins.
We adopt the Gaussian assumptions (37) and propose an iterative statistical linearization of the nonlinear system around a nominal trajectory. In this way, each iteration produces a LTV system from the linearization. We then perform the P-GVIMP algorithm 1 on the obtained linearized LTV system. The results are used in the next iteration for the SLR linearization. This process is performed until the linearization loop converges.
@@
Assuming the closed-loop discrete nonlinear stochastic dynamics at time $t=t_{i}$ as $X_{i+1}=f_{\rm cl}(X_{i})$, then the goal of SLR is to find a linear model $(A_{i},a_{i})$
$\hat{X}_{i+1}=A_{i}X_{i}+a_{i}$
to minimize the statistical error, i.e.,
@@
Taking the first-order derivative with respect to $a_{i}$ and let it be zero, we arrive at
@@
$\left(f_{\rm cl}(X_{i})-\mathbb{E}\left[f_{\rm cl}(X_{i})\right]-A_{i}\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)\right)\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)^{T},$
letting which be zero leads to
@@
From (39) and (40), the computation of $(a_{i}^{*},A_{i}^{*})$ involves the computation of the expectations of a nonlinear function $\mathbb{E}\left[f_{\rm cl}(X_{i})\right]$, which in this paper we use the sparse GH-quadrature rules mentioned in Lemma 1. Specifically, at time $t_{i}$, for a given desired approximation precision, $k_{q}$, we compute the set of weights and sigma points $\{(W^{i}_{l},\xi^{i}_{l})\}_{l=1}^{n^{k_{q}}}$ for the sparse quadrature, and perform the quadrature approximation
$\mathbb{E}\left[X_{i+1}\right]=\mathbb{E}\left[f_{\rm cl}(X_{i})\right]\approx\sum_{l=1}^{n^{k_{q}}}W_{l}f(\sqrt{P_{xx}}\;\xi_{l}+\mathbb{E}\left[X_{i}\right]),$
which is then substituted into (39) and (40). This procedure is iteratively performed forward in time. The iterative Parallel GVIMP is summarized in Algorithm 2.
7 experiments
This section presents the numerical experiment results for the proposed method. The experiments are aimed to answer the following questions:
  • 1.
    Q1: How much efficiency do we gain by leveraging the parallel computations on GPUs, compared with serial computations on CPUs?
  • 2.
    Q2: The effectiveness of the KL-proximal algorithm in high-DOF robot motion planning tasks, and its comparison with the GPMP2 baseline;
  • 3.
    Q3: The effectiveness of the iterative PGVIMP algorithm in motion planning tasks for nonlinear dynamical systems.
To answer these questions, we first demonstrate the efficiency improvement brought by the parallel computing paradigm on a laptop GPU. Secondly, we show the planning results for two $7$-DOF linear robotics kinematic problems. Thirdly, we show the planning results for a linearized system and a full nonlinear planar quadrotor system. Finally, we present some ablation study results and comparisons with baseline methods.

### v4

@@ 4.1 Sparse Motion Planning Factor Graph
which breaks down into calculating the expectations of $\psi$
@@
The motion planning factor graph is shown in the following Figure 1. It contains two types of factors: the $f_{i,i+1}(x_{i},x_{i+1})$ represents the prior factor enforcing the kinodynamics constraints, and the $f_{i}(x_{i})$ are the collision factors that encourage the trajectory to remain within obstacle-free regions.
@@
This factorized structure is leveraged to distribute the computation of $\psi_{d}$ on a GPU. Given the factor graph structure, we factorize the collision costs $\psi_{d}=\lVert h(x)\rVert^{2}_{\Sigma_{\rm obs}}$ into
@@
where the factorized potentials are defined as the collision cost on each state on the trajectory $\psi^{d}_{i}\triangleq\lVert h_{i}(q_{i})\rVert_{\sigma_{\rm obs}}^{i}$. The marginal Gaussian variables are linearly mapped from the joint Gaussian
@@ 4.2 Marginal updates
By Stein’s Lemma, (17a) and (17b) are written as
@@
Since $\psi(\mathbf{x})$ can be decomposed at the marginal level as
@@
where $\mathbf{x}_{i}=M_{i}\mathbf{x}$ is a subset of the variables relevant to the $i$th factor as defined in (19). Applying (21) to (20a), we obtain
@@
For the $i$-th factor, the gradient and expectation simplify to depend only on $\mathbf{x}_{i}$, since $\psi_{i}$ is a function of $\mathbf{x}_{i}$ alone. A similar result holds for (20b)
@@
Applying Stein’s Lemma again to (22) and (23) gives
@@ 4.3 Sparse Quadratures for Expectation Estimations
Our primary goal is to estimate the expectation (24) on the marginal levels. Gauss-Hermite quadrature approximations are widely used in filtering literature [33]. They are accurate and efficient for lower-dimensional problems. However, a vanilla tensor product of the full-grid sigma points scales exponentially with the dimension, making it computationally infeasible for higher-dimensional problems [48].
Sparse-grid quadrature rules (Smolyak’s rules) [49] ignore the cross terms among different dimensions to mitigate the issue, leading to polynomial dimensional dependence. This study adopts Smolyak’s rules in expectation estimations. To integrate a function $\varphi(\mathbf{x})$
$\int\varphi(\mathbf{x})\mathcal{N}(\mathbf{x};m,P)d\mathbf{x},$
@@
    Compute the $p$ roots of a pthp\;th order Hermite polynomial, also denoted as sigma points: $\xi=[\xi_{1},\dots,\xi_{p}].$
@@
    $\int\varphi(\mathbf{x})\mathcal{N}(\mathbf{x};m,P)d\mathbf{x}\approx\sum_{i}W_{i}\varphi(\sqrt{P}\xi_{i}+m).$ (25)
Lemma 1 (Smolyak’s rule complexity).
For an $n$-variate function, the number of computations needed by a sparse grid quadrature that is exact for $p=2k_{q}-1$ total polynomial order is bounded by [49]
@@
Algorithm 1 P-GVIMP (LTV system).
@@ 5  Perform P-GVIMP (Algorithm 1) on the linearized system
[…]
@@
Algorithm 2 i-P-GVIMP (Nonlinear system).
5 Belief Propagation and GPU-Parallel Collision Checking on Factor Graphs
We introduce the two key components that greatly increase the efficiency of the P-GVIMP algorithm for LTV systems. To begin with, we first state our main assumption on the number of discritizations and the state dimensions.
Assumption 1.
The discretization dimension $N$ (over $1000$ for dense trajectories) is far greater than the state space dimension $n$ (usually less than $20$ for robotics systems), i.e., $N\gg n$.
@@ 5.1 Belief propagation for marginal covariances update
Computing the marginal updates (22) and (23) requires computing the marginal mean and covariance of the Gaussian distributions $q_{i}$. A major bottleneck is computing the marginal covariances. A brute force inverse computation of the joint covariance from $\Sigma^{-1}_{\theta}$ has cubic complexity $O((N\times n)^{3})$.
Leveraging the sparse factor graph, this section introduces the Gaussian Belief Propagation (GBP) [28, 50] to compute the marginals over the factor graph. We write the Gaussian distribution $q\sim\mathcal{N}(\mu_{\theta},\Sigma_{\theta})$ in its canonical form $q(x)\propto\exp(-\frac{1}{2}x^{\top}\Lambda_{\theta}x+\eta_{\theta}^{\top}x)$ where $\Lambda_{\theta}=\Sigma_{\theta}^{-1}$ denotes the precision matrix, and $\eta_{\theta}=\Sigma_{\theta}^{-1}\mu_{\theta}$ denotes the information vector.
To simplify the calculation, assume a shifted Gaussian distribution $r\sim\mathcal{N}(0,\Sigma_{\theta})$, since only the marginal covariance is desired. Then $r(x)$ is written as
@@
The precision matrix $\Lambda_{\theta}$ has the following sparsity pattern
$\Lambda_{\theta}\!=\!\begin{bmatrix}\Lambda_{00}\!\!&\Lambda_{01}\!\!&&&&\\ \Lambda_{10}\!\!&\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &\Lambda_{11}\!\!&\Lambda_{12}\!\!&&&\\ &&&\dots&&\\ &&&&\!\!\Lambda_{(N-1)(N-1)}\!\!&\Lambda_{(N-1)N}\\ &&&&\!\!\Lambda_{N(N-1)}\!\!&\Lambda_{NN}\end{bmatrix}$,
and the Gaussian distribution (26) is factorized into $r(x)\propto\prod_{i=0}^{N}f_{i}(x_{i})\ \prod_{i=0}^{N-1}f_{i}(x_{i},x_{i+1})$ with
@@
Gaussian Belief Propagation is an algorithm for calculating the marginals of a joint distribution via local message passing between nodes in a factor graph [50]. Message passing on the factor graph falls into two types. The message passed from variables to factors, denoted as $m_{x_{i}\rightarrow f_{j}}$, and the message passed from factors to variables, denoted as $m_{f_{j}\rightarrow x_{i}}$. Messages are obtained by
@@
After computing all the messages in the factor graph, the beliefs of variables are obtained by taking the product of incoming messages: $b_{i}\left(x_{i}\right)=\prod_{s\in N(i)}m_{f_{s}\rightarrow x_{i}}$. Since the messages here are also Gaussian, we obtain the belief parameters $\Lambda_{bi}$ by
@@
The time complexity of inverting an $n\times n$ matrix is $O(n^{3})$, leading to total complexity $O(N\times n^{3})$ for GBP.
@@ 5.2 Parallel Collision Checking on Dense Trajectory
One of the most computationally heavy modules for motion planning algorithms is the collision-checking module on a dense trajectory [4, 12]. This is also the case in our framework, where the collision-checking is required in a probabilistic sense, i.e., the goal is to compute the expected collision costs with respect to the proposal distribution.
We compute the expected collision costs by distributing the computation to marginals in the factor graph (c.f., Section 4.2). At marginal levels, we deploy the sparse GH quadratures to compute the marginal collision costs (c.f., Section 4.3). The expectation approximation is done by computing (25). Thus, we distribute the expectation approximation by distributing the matrix multiplication (25) on the GPU. The graphical demonstration of the parallel collision checking is shown in Figure 2.
@@
5.3 The P-GVIMP Algorithm and Complexity Analysis
The P-GVIMP algorithm is summarized in the Algorithm 1. It serves as a sub-routine in planning for nonlinear systems.
Complexity Analysis. Denote the state dimension as $n$, and the time discretization number as $N+1$. The factor graph has $L=2N+3$ factors, the maximum dimension being $d=2n$. The main computational burdens are: (1) computing the nonlinear factors; and (2) computing marginal covariances.
Serial computation of the nonlinear factors has linear dependence on $N$, with total complexity $O(N\times n^{k_{q}})$, where $k_{q}$ is the desired precision in the GH-quadrature [16]. The parallel nonlinear factor computation on a GPU reduces the complexity to $O(n^{k_{q}})$. Experimentally, to obtain a precise estimate of the collision costs, $k_{q}\geq 6$ suffices. On the other hand, the complexity of computing the marginal covariances using GBP on the tree-structured factor graph is $O(N\times n^{3})$, which becomes the new bottleneck. The total algorithm complexity changes from $O(N\times n^{k_{q}})$ to $O(N\times n^{3})$.
@@ 6 Iterative P-GVIMP for Nonlinear Systems
Real robot dynamics are nonlinear. The P-GVIMP algorithm described in the previous sections is constrained in LTV systems. This section introduces our proposed Iterative P-GVIMP Algorithm (i-P-GVIMP) to handle nonlinearity.
@@ 6.1 Gaussian Assumptions for Nonlinear Transition
The Gaussian assumption is widely used in robotics applications [33, 32]. It assumes the state distribution follows
@@
where $\mathbb{E}_{\mathcal{N}}\left[\psi(\cdot)\right]$ denotes expectations under Gaussian distributions. We adopt the Gaussian assumptions (29) and propose an iterative statistical linearization of the nonlinear system around a nominal trajectory. The P-GVIMP Algorithm 1 is then performed on the obtained linearized LTV system. The results are used in the next iteration for the linearization. This process is performed until convergence.
@@
Assuming the closed-loop nonlinear stochastic dynamics at time $t=t_{i}$ as $X_{i+1}=f_{\rm cl}(X_{i})$, then SLR finds a linear model $\hat{X}_{i+1}=A_{i}X_{i}+a_{i}$ for $(A_{i},a_{i})$ to minimize the statistical error
@@
Taking the first-order derivative of (30) with respect to $a_{i}$ and letting the derivative be zero, we arrive at
@@
$\left(f_{\rm cl}(X_{i})-\mathbb{E}\left[f_{\rm cl}(X_{i})\right]-A_{i}\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)\right)\left(X_{i}-\mathbb{E}\left[X_{i}\right]\right)^{T}.$
Letting the above be zero leads to
@@
In (31) and (32), we use the sparse GH-quadrature rules in Lemma 1 to compute the expectations of a nonlinear function. At time $t_{i}$, for a given approximation precision, $k_{q}$, we compute the weights and sigma points $\{(W^{i}_{l},\xi^{i}_{l})\}_{l=1}^{n^{k_{q}}}$, and perform the quadrature approximation in (31) and (32)
$\mathbb{E}\left[X_{i+1}\right]=\mathbb{E}\left[f_{\rm cl}(X_{i})\right]\approx\sum_{l=1}^{n^{k_{q}}}W_{l}f(\sqrt{P_{xx}}\;\xi_{l}+\mathbb{E}\left[X_{i}\right]).$
This procedure is iteratively performed forward in time. The iterative P-GVIMP is summarized in Algorithm 2. Our proposed paradigm has profound connections with the previous works [15, 16], where we targeted to solve the efficiency issues and for nonlinear systems. Table 1 summarizes the connections and differences between P-GVIMP and GVIMP.
GVIMP P-GVIMP
Paradigm Planning-as-inference
Structure Sparse Planning Factor Graph
System Linear Nonlinear
Algorithm NGD SLR + KL-Proximal
Complexity $O(N\times n^{k_{q}})$ $O(N\times n^{3})$
Hardware CPU GPU
Table 1: Comparison between the proposed paradigm and the previous works on GVIMP [15, 16].
7 Numerical Experiments
This section presents the numerical experiment results for the proposed method. The experiments are aimed at understanding:
  • Q1: Efficiency gains by parallel computation on GPUs versus serial computation on CPUs;
  • Q2: The effectiveness of the KL-proximal algorithm in high-DOF robot motion planning tasks compared GPMP2;
  • Q3: The effectiveness of the iterative P-GVIMP algorithm in motion planning tasks for nonlinear dynamical systems.
This section answers these questions sequentially and concludes with ablation and comparison studies.